Emssj:

Mr.Baker is driving 100 miles . His average speed to the nearest 2 miles per hour is 56 miles per hour. What is the range of possible values for T the total time of the drive in hours

surjithayer:

distance=100 miles max. average speed=56+2=58 mph \[time~taken=\frac{ 100 }{ 58 }=\frac{ 50 }{ 29 }=1\frac{ 21 }{ 29 }~hrs.\] min.average speed=56-2=54 mph \[time~taken=\frac{ 100 }{ 54 }=\frac{ 50 }{ 27 }=1\frac{ 23 }{ 27 }~hrs\] \[T \in (1\frac{ 21 }{ 29 }hrs ,1\frac{ 23 }{ 27}hrs)~or~1\frac{ 21 }{ 29}hrs<T<1\frac{ 23 }{ 27 }hrs\]

surjithayer:

statement is not clear ,i thought it varies 2 mph either way of 56 mph.